a motorcycle can be driven 110 miles in 3 hours, how long will it take to travel 125 miles?
Answer by TimothyLamb(4379)

s = d / t
s = 110/3
s = 36.67 mph
---
t = d / s
t = 125/(110/3)
t = 3.4090909 hrs
---
answer:
3 hours 24 minutes 32.7 seconds
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php